OpenStudy (anonymous):

The quadratic formula. Classify statement as either true or false. t2 + 10 = 6t

OpenStudy (mimi_x3):

t^2 - 6t + 10 = 0 a = 1 , b = -6 , c = 10 Then substitute it to the quadratic formula

OpenStudy (chaise):

t^2 + 10 = 6t t^2 -6t + 10 = 0 b^2 - 4ac < 0; therefore false. The graph t^2 - 6t + 10 does not cut the x axis.